Composition of function:
composition of function is the nesting of one function into another function as an input of that function.
Let f(x) be the first function and g(x) be the second function So.
f(x)=x+8 and
g(x)=5x-3
nesting of g(x) into f(x) is
f(g(x))=?
it means the value of x for f(x) is equal to g(x).
f(g(x))=(5x-3)+8

Found 2 solutions by ikleyn, Theo:
Answer by ikleyn(52793)

.

Your logic, your reasoning and your writing are absolutely correct.


You even can go one step further and conclude then that  


    f(g(x)) = (5x-3)+8 = 5x+5 = 5*(x+1).



You can accept any of these equivalent forms as your final answer.


With only one notice from my side.

    When you consider the composition f(g(x)), it is more logical to call 
    g(x) as a first function, and call f(x) as a second function.


    It is not a tragic mistake to call them in different way, 
    but then you can easily confuse yourself and people around.


    But if you do not afraid to confuse yourself and people around,
    then you can call them in either way.



Answer by Theo(13342)

f(x) = x + 8
g(x) = 5x - 3
f(g(x)) = (5x - 3) + 8
that is correct.
what is the question?